OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −6

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

@tommylee21 try it fist.......

OpenStudy (ivettef365):

multiply the first equation by 2 and the second equation by -3 then add both equation

OpenStudy (anonymous):

so y = -3 as i added them both and they where 0 (3*2=6 2*-3 =-6 ) is that correct

OpenStudy (ivettef365):

2 (3x+2y=6) then you have 6x+4y = 12 -3(2x+3y=-6) then you have -6x-9y = 18 --------------- now subtract

OpenStudy (ivettef365):

I meant add

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

?

OpenStudy (primeralph):

looking good. Please carry on.

OpenStudy (anonymous):

so y =-6

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

thank u